OpenStudy (anonymous):

question: Your classmate solved (x+5)(x-1)=2 as follows: x+5=2 x-1=2 x=2-5 x=2+1 x= -3 x=3 Your teacher said this was wrong because the values did not check. Explain what is wrong with your classmate’s solution.

OpenStudy (amoodarya):

when x+5=2 and (x+5)(x-1)=2 so x=2-5 =-3 then x-1=-4 so (x+5)(x-1)=(-3+5)(-3-1)=(+2)(-4)=-8 #2

OpenStudy (anonymous):

ok thanks